Only c and e (1 point) On a typical day at a local popular mall, the number of shoplifters caught by mall security fluctuates from day to
Answer #1

c) Mean is 8.3 per day

We need to find x from 7 to 13 inclusive

P(7<I<13) = P(7) + P(8) + P(9) + P(10) + P(11) + P(12) + P(13)

P(7<< < 13) = e-8.3 8.39 7! T8! T9! 8.38 SA + + 8.310 8.311 8.312 8.313 10! 11 + 101+79) 0.677956

So, 0.678 (rounded to the 4 decimal places) is the answer.

e) mean = 8.3 per day

total opening time = 11 hours

So, mean per hour = 8.3/11 = 0.755 (rounded to 3 decimal places)

var = 0.755 (as in Poisson mean = var)

Please comment if any doubt. Thank you.